# PSN Game: Game-theoretic Prediction and Planning via a Player Selection Network

### v2

@@ Abstract.
While game-theoretic planning frameworks are effective at modeling multi-agent interactions, they require solving large optimization problems where the number of variables increases with the number of agents, resulting in long computation times that limit their use in large-scale, real-time systems. To address this issue, we propose i) PSN Game—a learning-based, game-theoretic prediction and planning framework that reduces runtime by learning a Player Selection Network (PSN); and ii) a Goal Inference Network (GIN) that makes it possible to use the PSN in incomplete information games where agents’ intentions are unknown. A PSN outputs a player selection mask that distinguishes influential players from less relevant ones, enabling the ego player to solve a smaller, masked game involving only selected players. By reducing the number of players in the game, and therefore reducing the number of variables in the corresponding optimization problem, PSN directly lowers computation time. The PSN Game framework is more flexible than existing player selection methods as it i) relies solely on observations of players’ past trajectories, without requiring full state, action, or other game-specific information; and ii) requires no online parameter tuning. Experiments in both simulated scenarios and human trajectory datasets demonstrate that PSNs outperform baseline selection methods in i) prediction accuracy; and ii) planning safety. PSNs also generalize effectively to real-world scenarios in which agents’ objectives are unknown without fine-tuning. By selecting only the most relevant players for decision-making, PSN Game offers a general mechanism for reducing planning complexity that can be seamlessly integrated into existing multi-agent planning frameworks.
Key words and phrases:
Multi-agent systems, Key player selection, Game-theoretic planning
1. Introduction
@@
Game-theoretic planning frameworks are widely used in robotics to model multi-agent interactions in terms of Nash (or other) equilibrium strategies, which can be identified by optimization-based algorithms Fridovich-Keil ([n.d.]); [”Sun”] (2025). Most existing approaches, however, focus on static scenarios or involve only a small number of agents (e.g., $\leq 5$), where computational efficiency is rarely a concern. In contrast, scenarios involving many dynamic agents demand frequent replanning, making computational efficiency a critical bottleneck. In such settings, the number of optimization variables, typically proportional to the number of agents, can grow rapidly and lead to significant delays. Prior work Laine et al. (2023); Fridovich-Keil et al. (2020) has shown that even in simple linear-quadratic games, computation time scales cubically with the total number of state and control variables for all agents, rendering game-theoretic planners impractical for large-scale use.
@@
Interestingly, similar limitations are observed in human driving. Prior research on drivers’ attention Gershon et al. (2019); Fang et al. (2021); Morando et al. (2018); Li et al. (2019); Xia et al. (2020) reveals that in dense traffic, monitoring all surrounding vehicles simultaneously is hard for human drivers, leading to distraction, delayed reactions, and increased accident risk. Although robots are not subject to distraction, they still face computational limitations that require longer solving times as the number of agents increases. In contrast, human drivers instinctively prioritize nearby vehicles that pose an immediate threat, such as those cutting in, merging, or failing to yield. This behavior suggests two key insights: i) it is infeasible to consider all agents at once; and ii) focusing on a strategically selected subset of agents is often sufficient for safe driving.
@@
Figure 1. Overview of our game-theoretic prediction and planning framework via the Player Selection Network (PSN) and the Goal Inference Network (GIN). At each timestep, the ego player (blue) observes other agents’ past trajectories and inputs them to PSN. The network selects important players (red) and excludes less relevant ones (green). The ego player then solves a masked game over the selected subset to obtain the Nash strategy $u_{k|k}^{i*}$ and updates its state using (1b).
[…]
To address these limitations, we propose PSN Game—a novel game-theoretic framework that learns a Player Selection Network (PSN) to reduce the computation time of game-solving for multi-agent trajectory prediction and planning. The PSN takes the interested agent and the surrounding agents’ past trajectories as input, and outputs a player selection mask identifying the most influential agents. Furthermore, we build a Goal Inference Network (GIN) that infers other agents’ objectives even when they are a priori unknown, making player selection practical even in incomplete information settings. Ultimately, the PSN, together with the GIN, constructs a smaller-scale masked game over the selected subset of agents, whose equilibrium solution can encode trajectory predictions and planned interactions.
(Fig. 1). Our contributions are fourfold:
@@
    We introduce an unsupervised Player Selection Network (PSN), trained with a differentiable dynamic game solver that enables backpropagation through the selection mask.
@@
    We empirically validate PSN Game in multi-agent simulations and real-world human trajectory datasets. PSN Game effectively reduces the scale of the game by 50% to 75%, resulting in a vast speedup in runtime compared to solving the full-player games.
PSN Game offers two key advantages over prior selection methods. i) Flexible data usage: at runtime, the PSN relies only on past trajectory data—either full-state (e.g., position and velocity) or partial-state (e.g., position only)—without requiring control inputs, cost functions, or other game-specific parameters. It also eliminates the need for online parameter tuning. ii) State-of-the-art performances: by leveraging spatio-temporal patterns in past trajectories, PSN achieves state-of-the-art performances in prediction and planning metrics over existing baseline selection methods. These features make PSN Game broadly applicable across diverse multi-agent planning and prediction scenarios.
2. Related Work
2.1. Game-theoretic Planning and Time Complexity Analysis
Game-theoretic planning has been widely applied to model diverse multi-agent interactions, including intention inference Peters et al. (2023); Liu et al. (2024); Le Cleac’h et al. (2021); Mehr et al. (2023), leadership inference Khan and Fridovich-Keil (2024); Hu et al. (2024), handling occluded agents in autonomous driving Zhang and Fisac (2021); Qiu and Fridovich-Keil (2024); Gupta and Fridovich-Keil (2023), and competitive scenarios such as racing Schwarting et al. (2021); Wang et al. (2021). These methods typically consider small-scale environments with relatively few agents, where real-time equilibrium computation using solvers such as Fridovich-Keil et al. (2020); Laine et al. (2023); Dirkse and Ferris (1995) is feasible. These Newton-style solvers work by iteratively approximating the underlying noncooperative game with a sequence of quadratic problems that can be solved analytically; however, the complexity of each subproblem scales cubically with the number of players, as illustrated in Fig. 2, which makes it nontrivial to solve larger-scale problems involving many agents.
To the best of our knowledge, absent any additional structure, e.g., homogeneity of agents’ objectives that admits a mean field representation Lasry and Lions (2007), this complexity presents a fundamental obstacle for deploying game-theoretic planning in large-scale, real-time multi-agent systems. Given that the horizon $T$ and state dimension $n$ for each agent are often fixed by the task, reducing the number of agents in the game remains the most viable strategy for improving computational performance.
Figure 2. Computation time using Fridovich-Keil ([n.d.]) vs. the number of players in the game given in Section 4.1.1. Time consumption grows cubically as player number increases.
[…]
2.2. Player Selection in Multi-agent Interactions
@@
Threshold-based methods highlight agents that breach a predefined distance threshold around the ego agent. This idea has been widely adopted in works such as Trautman and Krause (2010); Chen et al. (2017), where a safety margin is first defined to select key agents before applying prediction or planning frameworks. However, these methods heavily rely on threshold tuning, making them hard to generalize across scenarios.
Ranking-based methods, by contrast, aim to select the top-$k$ most influential agents according to certain heuristics, such as distance from the ego agent, Van den Berg et al. (2008); Snape et al. (2011) where the top $k$ nearest neighbors (kNNs) are identified. In learning-based approaches Everett et al. (2018); Chen et al. (2019); Wang et al. (2025), pooling mechanisms are employed to fix the input size to a neural network trained to identify nearest neighbors. In the context of game-theoretic planning, Chahine et al. (2023) introduced ranking metrics based on, e.g., the sensitivity of the ego agent’s cost function to other agents’ state.
@@
In contrast, our proposed PSN Game addresses the limitations of both threshold-based and ranking-based approaches. PSN Game is lightweight in information requirements, relying only on past position observations at runtime without needing access to potentially privileged information (e.g., agents’ velocity, control inputs, player cost functions, or other game-specific information). In particular, although cost functions are used during training to generate trajectories, PSN Game does not require that information at test time, enabling flexible deployment across different environments with model mismatch. Experiments demonstrate that it not only improves computational efficiency but also enhances ego-agent trajectory smoothness and length, while maintaining safety performance across a range of multi-agent navigation scenarios.
3. Learning to Identify Key Players in Noncooperative Interactions
3.1. Preliminaries: Nash Games
A finite horizon, discrete-time, open-loop Nash game with $N$ agents is characterized by their respective states $x_{k}^{i}\in\mathbb{R}^{n}$} and control inputs $u_{k}^{i}\in\mathbb{R}^{m}$}, $i\in\{1,\dots,N\}\equiv[N]$}, over time $k\in[T]$, where $T$ is the planning horizon. Each agent $i$’s state transition from time $k$ to $k+1$ is governed by the dynamics $x_{k+1}^{i}=f^{i}(x_{k}^{i},u_{k}^{i})$}. We denote agent $i$’s trajectory and control sequence compactly as $\textbf{x}^{i}\coloneqq(x_{0:T}^{i})$} and $\textbf{u}^{i}\coloneqq(u_{0:T-1}^{i})$}; and similarly, all agents’ states/control input at time $k$ by $\textbf{x}_{k}\coloneqq(x_{k}^{1:N})$, $\textbf{u}_{k}\coloneqq(u_{k}^{1:N})$ and their trajectory/control sequence over time by $\textbf{x}\coloneqq(\textbf{x}_{0:T})$, $\textbf{u}\coloneqq(\textbf{u}_{0:T-1})$, respectively. The function $J^{i}(\textbf{x},\textbf{u};\theta^{i})\coloneqq\sum_{k=0}^{T}c_{k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$ defines the $i^{\text{th}}$} agent’s cumulative cost, with game-specific parameter $\theta^{i}$. The game is thus fully characterized by all agents’ costs parameterized by $\theta\coloneqq(\theta^{1:N})$, initial states $\textbf{x}_{0}$, and dynamics $\textbf{f}\coloneqq(f^{1:N})$, and is denoted as $\Gamma(\textbf{x}_{0},\textbf{f};\theta)$}. In this game, each agent aims to minimize its cost while adhering to dynamics, i.e., agent $i$ solves the optimization problem minimizing the time-cumulative sum of private cost $c_{p,k}^{i}$ and shared cost $c_{s,k}^{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}c_{k}^{ij}$, where $c_{k}^{ij}$ is the mutual shared cost:
@@ 3.1. Preliminaries: Nash Games
which is inherently coupled with each other agent’s ($j\neq i$) own problem.
@@
concerning state and control trajectories of agent $i$ ($\textbf{x}^{i},\textbf{u}^{i}$) and other agents ($\textbf{x}^{-i},\textbf{u}^{-i}$) are satisfied for all $\textbf{x}^{i},\textbf{u}^{i}$ that remain feasible with respect to (1b), then $\textbf{u}^{i*}$ is called an Open-Loop Nash equilibrium (OLNE) strategy with $\textbf{x}^{i*}$ being the corresponding OLNE state trajectory for agent $i$. This inequality indicates that no agent can reduce their cost by unilaterally deviating from $\textbf{u}^{i*}$} Başar and Olsder (1998).
Differentiability of Nash game solvers. A key property of solutions $(\textbf{x}^{*},\textbf{u}^{*})$ to generalized Nash problems is that they are (directionally) differentiable with respect to problem parameters $\theta$, as discussed in Liu et al. (2023). In our work, we employ the solver in [”Sun”] (2025), which leverages implicit differentiation to efficiently compute those derivatives and thus enables backpropagation during training.
3.2. Masked Nash Game with Selected Players
@@
Given an open-loop Nash game $\Gamma(\textbf{x}_{0},\textbf{f};\theta)$}, a masked Nash game for agent $i$ is denoted as ${\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta})}$, where we define the set of unmasked agents (agents that are included in the game) $U^{i}\coloneqq\{j\neq i\penalty 10000\ |\penalty 10000\ m^{ij}=1\}$} and let
$\tilde{\theta}\coloneqq(\theta^{j})_{j\in U^{i}},\quad\tilde{\textbf{x}}_{k}$ $\coloneqq(x_{k}^{j})_{j\in U^{i}},\quad\tilde{\textbf{f}}\coloneqq(f^{j})_{j\in U^{i}},$ (4)
$\tilde{c}_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$ $\coloneqq\sum_{j\in U^{i}}c_{s,k}^{ij}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i}).$
The masked game preserves only the specific parameters and states corresponding to those agents most relevant to agent $i$. Agent $i$ then solves the masked Nash game $\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta},M^{i})$}, and employs its masked OLNE strategy $\tilde{\textbf{u}}^{i*}$} as its own plan.
3.3. Player Selection Network (PSN)
Given this construction, we now turn to the challenge of identifying the set of key players that ego agent $i$ should consider in its masked game model of the interaction in order to maintain safety and efficiency while minimizing the computational burden of solving the corresponding game to obtain its strategy at each time $k$. We propose the Player Selection Network (PSN), which is a neural network whose task is to infer a mask $M^{i}$ which balances performance with sparsity. Based on the available information, we present two variants of the network: PSN-Full, whose input is all agents’ past states $\textbf{x}_{0:K}$, and PSN-Partial [^1] While PSN-Partial is capable of selecting key agents from partial observations, we emphasize that solving the masked game still requires the knowledge of all agents’ full states. Prior works Peters et al. (2021); Qiu and Fridovich-Keil (2024) investigate methods to estimate full states of agents from partial observations via inverse games. PSN-Partial is designed to integrate with methods where only partial states of agents are needed., whose input is only a partial observation of state, i.e. $\{h(\textbf{x}_{k})\}_{k=0}^{K}$ for a known function $h$. For example, $h(\textbf{x}_{k})$ could return only the Cartesian position of all agents on the road, and not their velocities. Both variants’ output is the inferred mask $M^{i}$. Player $i$ will then solve the masked game $\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta},M^{i})$} to obtain his strategy $\tilde{\textbf{u}}_{K:K+T}^{i}$.

### v3

# A Player Selection Network for Scalable Game-Theoretic Prediction and Planning

@@ Abstract
While game-theoretic planning frameworks are effective at modeling multi-agent interactions, they require solving large optimization problems where the number of variables increases with the number of agents, resulting in long computation times that limit their use in large-scale, real-time systems. To address this issue, we propose i) PSN Game—a learning-based, game-theoretic prediction and planning framework that reduces game size by learning a Player Selection Network (PSN); and ii) a Goal Inference Network (GIN) that makes it possible to use the PSN in incomplete-information games where other agents’ intentions are unknown to the ego agent. A PSN outputs a player selection mask that distinguishes influential players from less relevant ones, enabling the ego player to solve a smaller, masked game involving only selected players. By reducing the number of players included in the game, PSN shrinks the corresponding optimization problems, leading to faster solve times. The PSN Game framework is more flexible than existing player selection methods as it i) relies solely on observations of players’ past trajectories, without requiring full state, action, or other game-specific information; and ii) requires no online parameter tuning. Experiments in both simulated scenarios and real-world pedestrian trajectory datasets show that PSN is competitive with, and often improves upon, the evaluated explicit game-theoretic selection baselines in i) prediction accuracy and ii) planning safety. Across scenarios, PSN typically selects substantially fewer players than are present in the full game, thereby reducing game size and planning complexity. PSN also generalizes to settings in which agents’ objectives are unknown, via the GIN, without test-time fine-tuning. By selecting only the most relevant players for decision-making, PSN Game provides a practical mechanism for reducing planning complexity that can be integrated into existing multi-agent planning frameworks.
1 Introduction
@@
Game-theoretic planning frameworks are widely used in robotics to model multi-agent interactions in terms of Nash (or other) equilibrium strategies, which can be identified by optimization-based algorithms (Fridovich-Keil et al., 2020; Sun, 2025). Most existing approaches, however, focus on static scenarios or involve only a small number of agents (e.g., $\leq 5$), where computational efficiency is rarely a concern. In contrast, scenarios involving many dynamic agents demand frequent replanning, making computational efficiency a critical bottleneck. In such settings, the number of optimization variables, typically proportional to the number of agents, can grow rapidly and lead to significant delays. Prior work (Laine et al., 2023; Fridovich-Keil et al., 2020) has shown that even in simple linear-quadratic games, computation time scales cubically with the total number of state and control variables for all agents, rendering game-theoretic planners impractical for large-scale use.
@@
Similar limitations are observed in human driving behaviors. Prior research on drivers’ attention (Gershon et al., 2019; Fang et al., 2021; Morando et al., 2018; Li et al., 2019; Xia et al., 2020) shows that in dense traffic, monitoring all surrounding vehicles simultaneously is difficult. Although robots are not subject to distraction, they still face onboard compute limits in repeated solver-in-the-loop planning as the number of agents increases. These observations suggest two key insights: i) it is infeasible to consider all agents at once; and ii) focusing on a strategically selected subset of agents is often sufficient for safe interaction.
@@
Figure 1: Overview of our game-theoretic prediction and planning framework via the Player Selection Network (PSN) and the Goal Inference Network (GIN). At each time step, the ego player (blue) observes other agents’ past trajectories and inputs them to PSN. This explicit reduced-game construction supports repeated receding-horizon solves under onboard compute limits. The network selects important players (red) and excludes less relevant ones (green). The ego player then solves a masked game over the selected subset to obtain the Nash strategy $u_{k|k}^{i*}$ and updates its state using equation 1b.
To address these limitations, we propose PSN Game—a novel game-theoretic framework that learns a Player Selection Network (PSN) to reduce game size in multi-agent trajectory prediction and planning. The PSN takes the ego agent and surrounding agents’ past trajectories as input and outputs a player-selection mask identifying the most influential agents. Furthermore, we build a Goal Inference Network (GIN) that infers other agents’ objectives even when they are unknown a priori, making player selection practical even in incomplete-information settings. Ultimately, the PSN, together with the GIN, constructs a smaller-scale masked game over the selected subset of agents, whose equilibrium solution can encode trajectory prediction and planning interactions (Fig. 1). Our contributions are fourfold:
@@
    We introduce an objective-driven Player Selection Network (PSN), trained with a differentiable dynamic game solver that enables backpropagation through the selection mask.
@@
    We empirically validate PSN Game in multi-agent simulations and real-world pedestrian trajectory datasets. PSN Game typically reduces the game size by 50% to 75% while maintaining strong prediction and planning performance relative to full-player games and the evaluated baselines.
PSN Game offers two key advantages over prior selection methods: i) Flexible data usage: at runtime, the PSN relies only on past trajectory data—either full-state (e.g., position and velocity) or partial-state (e.g., position only)—without requiring control inputs, cost functions, or other game-specific parameters. It also eliminates the need for online parameter tuning. ii) Strong empirical performance: by leveraging spatio-temporal patterns in past trajectories, PSN achieves strong and consistent performance in prediction and planning metrics relative to the evaluated explicit selection baselines. These features make PSN Game broadly applicable across diverse multi-agent planning and prediction scenarios.
2 Related Work
2.1 Game-theoretic Planning and Time Complexity Analysis
Game-theoretic planning has been widely applied to model diverse multi-agent scenarios, including intention inference (Le Cleac’h et al., 2021; Mehr et al., 2023; Peters et al., 2023; Liu et al., 2024b), hierarchical interactions (Khan and Fridovich-Keil, 2024; Hu et al., 2024; Khan et al., 2026), handling occluded agents in autonomous driving (Zhang and Fisac, 2021; Gupta and Fridovich-Keil, 2023; Qiu and Fridovich-Keil, 2024), and robot formation tasks (Liu et al., 2024a; Qiu et al., 2026). These methods typically consider small-scale environments with relatively few agents, where real-time equilibrium computation using solvers such as (Dirkse and Ferris, 1995; Fridovich-Keil et al., 2020; Laine et al., 2023) is feasible. These Newton-style solvers work by iteratively approximating the underlying noncooperative game with a sequence of quadratic problems that can be solved analytically; however, the complexity of each subproblem scales cubically with the number of players, as illustrated in Fig. 2, which makes it nontrivial to solve larger-scale problems involving many agents.
To the best of our knowledge, absent any additional structure, e.g., homogeneity of agents’ objectives that admits a mean field representation (Lasry and Lions, 2007), this complexity presents a fundamental obstacle for deploying game-theoretic planning in large-scale, real-time multi-agent systems. Given that the horizon $T$ and state dimension $n$ for each agent are often fixed by the task, reducing the number of agents in the game remains the most viable strategy for improving computational performance.
Figure 2: Computation time from (Fridovich-Keil, ) versus the number of players in the game given in Section 4.1.1. Time consumption grows cubically as player number increases.
2.2 Player Selection in Multi-agent Interactions
@@
Threshold-based methods highlight agents that breach a predefined distance threshold around the ego agent. This idea has been widely adopted in works such as (Trautman and Krause, 2010; Chen et al., 2017), where a safety margin is first defined to select key agents before applying prediction or planning frameworks. However, these methods heavily rely on threshold tuning, making them hard to generalize across scenarios.
Ranking-based methods, by contrast, select the top-$k$ most influential agents according to a heuristic, often by distance to the ego agent; common examples include nearest-neighbor selection (Van den Berg et al., 2008; Snape et al., 2011). In learning-based approaches (Everett et al., 2018; Chen et al., 2019; Wang et al., 2025), pooling mechanisms are employed to fix the input size to a neural network trained to identify nearest neighbors. In the context of game-theoretic planning, Chahine et al. (2023); Vakil and Pierson (2026) introduced ranking metrics based on, e.g., the sensitivity of the ego agent’s cost to other agents’ state.
@@
In contrast, our proposed PSN Game addresses the limitations of both threshold-based and ranking-based approaches. PSN Game is lightweight in information requirements at runtime. PSN-Partial uses only past position observations, while PSN-Full uses past state trajectories; neither variant requires controls, cost functions, or other game-specific information at test time. In particular, although cost functions are used during training to generate trajectories, PSN Game does not require that information at test time, enabling flexible deployment across different environments with model mismatch. Experiments show that PSN remains competitive on safety and often produces smoother, shorter ego trajectories relative to the evaluated baselines across a range of multi-agent navigation scenarios.
3 Learning to Identify Key Players in Noncooperative Interactions
3.1 Preliminaries: Nash Games
A finite horizon, discrete-time, open-loop Nash game with $N$ agents is characterized by their respective states $x_{k}^{i}\in\mathbb{R}^{n}$} and control inputs $u_{k}^{i}\in\mathbb{R}^{m}$}, $i\in\{1,\dots,N\}\equiv[N]$}, over time $k\in[T]$, where $T$ is the planning horizon. Each agent $i$’s state transition from time $k$ to $k+1$ is governed by the dynamics $x_{k+1}^{i}=f^{i}(x_{k}^{i},u_{k}^{i})$}. We denote agent $i$’s trajectory and control sequence compactly as $\textbf{x}^{i}\coloneqq(x_{0:T}^{i})$} and $\textbf{u}^{i}\coloneqq(u_{0:T-1}^{i})$}; and similarly, all agents’ states/control inputs at time $k$ by $\textbf{x}_{k}\coloneqq(x_{k}^{1:N})$, $\textbf{u}_{k}\coloneqq(u_{k}^{1:N})$ and their trajectory/control sequence over time by $\textbf{x}\coloneqq(\textbf{x}_{0:T})$, $\textbf{u}\coloneqq(\textbf{u}_{0:T-1})$, respectively. The function $J^{i}(\textbf{x},\textbf{u};\theta^{i})\coloneqq\sum_{k=0}^{T}c_{k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})$ defines the $i^{\text{th}}$} agent’s cumulative cost, with game-specific parameter $\theta^{i}$. The game is thus fully characterized by all agents’ costs parameterized by $\theta\coloneqq(\theta^{1:N})$, initial states $\textbf{x}_{0}$, and dynamics $\textbf{f}\coloneqq(f^{1:N})$, and is denoted as $\Gamma(\textbf{x}_{0},\textbf{f};\theta)$}. In this game, each agent aims to minimize its cost while adhering to dynamics, i.e., agent $i$ solves the optimization problem minimizing the time-cumulative sum of private cost $c_{p,k}^{i}$ and shared cost $c_{s,k}^{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}c_{k}^{ij}$, where $c_{k}^{ij}$ is the mutual shared cost:
@@ 3.1 Preliminaries: Nash Games
so each agent’s optimization problem is inherently coupled with the problems of the other agents.
@@
concerning state and control trajectories of agent $i$ ($\textbf{x}^{i},\textbf{u}^{i}$) and other agents ($\textbf{x}^{-i},\textbf{u}^{-i}$) are satisfied for all $\textbf{x}^{i},\textbf{u}^{i}$ that remain feasible with respect to equation 1b, then $\textbf{u}^{i*}$ is called an Open-Loop Nash equilibrium (OLNE) strategy with $\textbf{x}^{i*}$ being the corresponding OLNE state trajectory for agent $i$. This inequality indicates that no agent can reduce their cost by unilaterally deviating from $\textbf{u}^{i*}$} (Başar and Olsder, 1998).
Differentiability of Nash game solvers A key property of solutions $(\textbf{x}^{*},\textbf{u}^{*})$ to generalized Nash problems is that they are (directionally) differentiable with respect to problem parameters $\theta$, as discussed in (Liu et al., 2023). In our work, we employ the solver in (Sun, 2025), which leverages implicit differentiation to efficiently compute those derivatives and thus enables backpropagation during training.
3.2 Masked Nash Game with Selected Players
@@
Given an open-loop Nash game $\Gamma(\textbf{x}_{0},\textbf{f};\theta)$}, a masked Nash game for agent $i$ is denoted as ${\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta})}$, where we define the set of unmasked agents (agents that are included in the game) $U^{i}\coloneqq\{j\neq i~|~m^{ij}=1\}$} and let
$\tilde{\theta}\coloneqq(\theta^{j})_{j\in U^{i}},\quad\tilde{\textbf{x}}_{k}$ $\coloneqq(x_{k}^{j})_{j\in U^{i}},\quad\tilde{\textbf{f}}\coloneqq(f^{j})_{j\in U^{i}},\quad\tilde{c}_{s,k}^{i}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i})\coloneqq\sum_{j\in U^{i}}c_{s,k}^{ij}(\textbf{x}_{k},\textbf{u}_{k};\theta^{i}).$ (4)
The masked game retains only the states, dynamics, and parameters associated to those agents most relevant to agent $i$. Agent $i$ then solves the masked Nash game $\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta},M^{i})$}, and employs its masked OLNE strategy $\tilde{\textbf{u}}^{i*}$} as its own plan.
3.3 Player Selection Network (PSN)
Given this construction, we now turn to the challenge of identifying the set of key players that ego agent $i$ should consider in its masked game model of the interaction in order to maintain safety and efficiency while minimizing the computational burden of solving the corresponding game to obtain its strategy at each time $k$. We propose the Player Selection Network (PSN), which is a neural network whose task is to infer a mask $M^{i}$ that balances performance with sparsity. Based on the available information, we present two variants of the network: PSN-Full, whose input is all agents’ past states $\textbf{x}_{0:K}$, and PSN-Partial [^1] While PSN-Partial is capable of selecting key agents from partial observations, we emphasize that solving the masked game still requires the knowledge of all agents’ full states. Prior works (Peters et al., 2021; Qiu and Fridovich-Keil, 2024) investigate methods to estimate full states of agents from partial observations via inverse games. PSN-Partial is designed to integrate with methods where only partial states of agents are needed., whose input is only a partial observation of state, i.e. $\{h(\textbf{x}_{k})\}_{k=0}^{K}$ for a known function $h$. For example, $h(\textbf{x}_{k})$ could return only the Cartesian position of all agents on the road, and not their velocities. Both variants’ output is the inferred mask $M^{i}$. Player $i$ will then solve the masked game $\Gamma^{i}(\tilde{\textbf{x}}_{0},\tilde{\textbf{f}};\tilde{\theta},M^{i})$} to obtain its strategy $\tilde{\textbf{u}}_{K:K+T}^{i}$.